You draw in a deep breath on a chilly day, inhaling 3.6 L of 0∘C air. If the pressure in your lungs is a constant 1.0 atm, how much heat must your body supply to warm the air to your 37∘C internal body temperature? Assume that for air Cp = 29.1 J/(K⋅mol) as it consists mostly of nitrogen
Answer #1

PV = nRT

n = PV / RT

n = 101325 * 0.0036 / 8.314 * 273.15

n = 0.160623057 moles

so,

Q = 0.160623057 * 29.1 * 37

Q = 172.94 J